OpenStudy (sidsiddhartha):

stuck on this one: show that \[\Large \sum_{n=0}^{\infty}\frac{ z^n }{ 2^{n+1} }\] and \[\Large \sum_{n=0}^{\infty}\frac{ (z-i)^n }{ (2-i)^{n+1}}\] are analytic continuations of each other

OpenStudy (ikram002p):

ok i can show its analytic , but what does continuations of each other of each other mean?

OpenStudy (sidsiddhartha):

same here i'm doing this kinda prob. for the first time wiki says something about it-- http://en.wikipedia.org/wiki/Analytic_continuation

OpenStudy (sidsiddhartha):

i think we have to show that those power series represents a same function @ikram002p

OpenStudy (sidsiddhartha):

for this function \[\Large \sum_{n=0}^{\infty}\frac{ z^n }{ 2^{n+1} }\] i used ratio test and found out that this series converges for \( |z|<2\)  which is a geometric series with first term 1/2 and common ratio Z/2 so it will be=---- \[\frac{ 1/2 }{ 1-(z/2) }=1/(2-z)\] it is correct? @ikram002p